# Turbulent scaling law in Ogata Kōrin’s Red and White Plum Blossoms

### v1

@@ V Concluding discussion
We have analyzed the luminance of the flowing river of Red and White Plum Blossoms. The spectrum of the luminance with the windowed Fourier transform followed closely $k^{-5/3}$ in the two different regions of the river. This suggests that the luminance spectrum is consistent with the Obukhov–Corrsin spectrum of the passive scalar advected by HIT. We then calculate the even-order structure functions of the luminance up to the 6-th order and found that they showed the power-law behavior although its range was limited. The power-law exponents of the luminance structure functions were also approximately consistent with the exponents with the intermittent exponents of the passive scalar in turbulence. Therefore we conclude that the flowing-water pattern of Red and White Plum Blossoms follows the universal (intermittent) scaling law of the passive scalar in the turbulent flow.
@@ V.1 Robustness of the scaling
Is the quantitative agreement between the stylized artwork and the natural phenomenon real or not? Of course, it can be just an artifact of the image processing, the photographic lighting or other factors, which are not related to the painting itself. To address this issue, we take two other JPG images of Red and White Plum Blossoms, which are supposedly independent from the one shown in Fig. 1. We analyze them in exactly the same manner.
@@ V.2 Why consistent?
The last question would be why and how Kōrin created the flowing water pattern, or suiryu moyō in Japanese, resulting in approximately the same intermittent scaling law of the passive scalar turbulence.
@@
In the case of Kōrin, the same argument about his ability of observation applies without doubt. Regarding the materials used for Red and White Plum Blossoms, recent X-ray analyses MOA Museum of Art and  Tokyo National Research Institute for Cultural Properties (2005) (eds.); Uchida et al. discovered that gold is not used in the foreground or background of the flowing river and that silver is used all over the river. But what was used precisely there is not identified yet. It may be of some future use to estimate the characteristic length of the small scales observed in this study. From the luminance spectrum shown in Fig. 4, the end of the $k^{-5/3}$-like part is estimated as $k\simeq 15$ [(cm)-1]. It corresponds to $0.42$ [cm], which can be regarded as the Kolmogorov dissipation length for the case of unit Schmidt number or the Obukhov–Corrsin length for the case of the low Schmidt number.
@@
Regarding the technique of Kōrin used for the artwork, a recent studyUchida et al. involving present-day artists using similar materials argue that the flowing pattern is likely drawn freehand in one sitting and that use of a stencil could not explain the expressed sense of momentum (but use of a stencil is not completely ruled out). Paper marbling (suminagashi in Japanese) or water transfer printing is ruled out. Indeed, looking at Fig. 1 or enlarged photographs of various parts of the painting in Ref. MOA Museum of Art and  Tokyo National Research Institute for Cultural Properties (2005), eds., we can imagine that Kōrin drew without hesitation the lines of the flowing patterns. To have a feeling of the scale hierarchy Kōrin depicted, specifically in the panel number 98 on page 73 of Ref. MOA Museum of Art and  Tokyo National Research Institute for Cultural Properties (2005), eds., which corresponds to a partenl of the region S2 in this paper, we observe that in that panel the scale ratio of the largest line widththi to the smallest is about $85$. This makes a good room to reproduce the broken self-similarity of the passive scalar turbulence.
@@
Still the question remains: How did Kōrin then reproduce the intermittent scaling law in one sitting? Let us suppose that a creator of a stylized artwork discard many details of the object and picks up a few essential elements of it. Perhaps what Kōrin picked up from the complexity of the turbulent flow reflects what we now know as the universal scaling law of the passive scalar turbulence. Specifically his irregular hierarchy of the flow pattern ultimately ended up reproducing it. Further investigation of Kōrin’s construction of the flow may lead to a method that is in the same spirit of the recent mathematical construction of turbulent weak solutions to the Euler equations, known as the convex integration Buckmaster et al. (2019). This may be a far-fetched idea. But, if it succeeds, then we see an example of the provocation made in Ref. Warhaft, 2022: “Artists can complement scientific analysis and help in developing a comprehensive theory of turbulence”.
@@ Acknowledgements.
I thank the MOA Museum of Art, the Wikimedia Commons and the Google Art & Culture initiative for making the artwork available in the high-resolution JPG format. I gratefully acknowledge discussions with Kohei Okuyama. This work was supported by JSPS KAKENHI Grant Number JP19K03669 and Number JP23K03247.

### v2

@@ V Concluding discussion
We have analyzed the luminance of the flowing river and the bark of the white plum-tree trunk in Red and White Plum Blossoms. The spectrum of the luminance with the windowed Fourier transform followed closely $k^{-5/3}$ in the two different regions of the river and the bark region. This suggests that the luminance spectrum is consistent with the Obukhov–Corrsin spectrum of the passive scalar advected by HIT. To examine this consistency further, we calculate the even-order structure functions of the luminance up to the 6th order and found that they showed the power-law behavior although its range was limited. The power-law exponents of the luminance structure functions were also approximately consistent with the exponents with the intermittent exponents of the passive scalar in turbulence. Therefore we conclude that the flowing-water pattern and the depiction of the bark of the white plum-tree trunk of Red and White Plum Blossoms follow the universal (intermittent) scaling law of the passive scalar in the turbulent flow.
@@ V.1 Robustness of the scaling
Is the quantitative agreement between the stylized artwork and the natural phenomenon real or not? Of course, it can be an artifact in the process of making an electronic image such as the image processing, the photographic lighting and so forth, which are not related to the painting itself. To address this issue, we take two other JPG images of Red and White Plum Blossoms, which are supposedly independent from the one shown in Fig. 1. Here we limit ourselves to the river regions where the number of data points are relatively large. We analyze them in exactly the same manner.
@@ V.2 Why consistent?
The last question would be why and how Kōrin created the flowing water pattern, or suiryu moyō in Japanese, and the bark of the plum-tree trunk, resulting in approximately the same intermittent scaling law of the passive scalar turbulence.
@@
In the case of Kōrin, the same argument about his ability of observation applies without doubt. We first consider the bark of the trunk of the white plum tree (the region C3 and S3). The painting technique used for this part is well-known as tarashikomi in Japanese Lippit (2007); Carpenter (2012); MOA Museum of Art and  Tokyo National Research Institute for Cultural Properties (2005) (eds.), which is developed by the Kōrin school. In this technique, the overlaied paint is placed while the underlaied paint is still wet. In the bark part of Red and White Plum Blossoms, its texture and lichens on it are depicted with tarashikomiMOA Museum of Art and  Tokyo National Research Institute for Cultural Properties (2005) (eds.). Although the bark does not look like a passive scalar vigorously stirred by a turbulent flow, the tarashikomi technique can give rise to turbulence locally in the paints to some extent. As a result, we observed the turbulent scaling law there. This is an explanation we propose for the bark region. It is in line with the argument for the Bachelor scaling in The Starry NightMa et al. (2024). Still, how the turbulent stirring we propose, if it is true, reproduces the turbulent scaling law when the paints are dried remains to be understood. Now, for a possible future study, let us estimate the dissipation scale. From the luminance spectrum shown in Fig. 4, the end of the $k^{-5/3}$-like part is estimated as $k\simeq 15$ [(cm)-1]. It corresponds to $0.42$ [cm], which can be regarded as the Kolmogorov dissipation length for the case of unit Schmidt number or the Obukhov–Corrsin length for the case of the low Schmidt number.
@@
The same explanation may not be applicable to the flowing water pattern in the regions C1 and C2 or S1 and S2 since the painting technique used there is not known and a matter of debate. A recent studyUchida et al. involving present-day artists using similar materials argues that the flowing pattern is likely drawn freehand in one sitting and that use of a stencil could not explain the expressed sense of momentum (but use of a stencil is not completely ruled out). Paper marbling (suminagashi in Japanese) or water transfer printing is ruled out. Indeed, looking at Fig. 1 or enlarged photographs of various parts of the painting in Ref. MOA Museum of Art and  Tokyo National Research Institute for Cultural Properties (2005), eds., we can imagine that Kōrin drew without hesitation the lines of the flowing patterns. One possibility is that finishing in a single sitting with some momentum can induce an effect similar to tarashikomi, namely, local turbulent stirring, which leads to the scaling law we observed. Here we infer that the effect results in variations in ink density within the brown curves of the swirling pattern.
@@
Nevertheless, for the flowing water pattern in particular, there is another possibility. That is, the turbulent scaling law there is not due to the turbulent stirring of the painting material, but due to Kōrin’s depiction. Specifically, his irregular hierarchy of the flow pattern ultimately ended up reproducing it. This is a far fetched speculation, though. Further investigation of Kōrin’s construction of the flow with much higher resolutions will tell us whether it is likely. If it is, then his way of depiction may leads a method that is in the same spirit of the recent mathematical construction of turbulent weak solutions to the Euler equations, known as the convex integration Buckmaster et al. (2019). If it succeeds, then we see an example of the provocation made in Ref. Warhaft, 2022: “Artists can complement scientific analysis and help in developing a comprehensive theory of turbulence”.
As a final remark, there is still a possibility that the scaling law of the parts of Red and White Plum Blossoms has nothing to do with fluid turbulence. If this is the case, it could imply that there is a universality class of scalar fields having the same intermittent scaling law and that the passive scalar turbulence and the parts of Red and White Plum Blossoms belong to this universality class.
@@ Acknowledgements.
I thank the MOA Museum of Art, the Wikimedia Commons and the Google Art & Culture initiative for making the artwork available in the high-resolution JPG format. I gratefully acknowledge Kohei Okuyama for discussions, one referee for critical comments and the other referee for suggesting analysis of the bark on the trunk of the left plum tree (the regions C3 and S3). This work was supported by JSPS KAKENHI Grant Number JP19K03669 and Number JP23K03247.
@@
Appendix A Third-order structure function
In this appendix we consider the 3rd-order structure function of the luminance in the square regions, S1, S2 and S3. As we mentioned in Sec. IV, convergence of the odd-order moments involving cancellation requires more data points than the even-order ones do. Accordingly, the results presented here may not be sufficiently converged. Nevertheless, we plot $S_{3}(r)$ in Fig. 8(a). The figure shows that its sign varies: it is positive in S2 and S3, but it is not definite in S1. The wiggly behaviors of the curves for S1 and S2 indicate the lack of convergence, though. In the numerical simulation of the passive scalar turbulence, the skewness of the passive scalar increment is found to be a small positive constant over the inertial-convective range Watanabe and Gotoh (2004). However, since the forcing added to the passive scalar equation in Ref.Watanabe and Gotoh, 2004 is a large-scale random forcing, it is argued that the skewness should vanish, provided that the turbulence in the large scales is isotropic enoughWatanabe and Gotoh (2004). If the passive scalar is maintained by the mean scalar gradient, it is known that the skewness is not zero Warhaft (2000). We do not pursue our comparison in his direction however.
The most important 3rd-order moment in the passive scalar turbulence is the Yaglom $4/3$ law that involves the velocity increment. For the painting, if there is some object that can be interpreted as the velocity field, then we can study whether the Yaglom $4/3$ lawYaglom (1949); Warhaft (2000); Gotoh and Yeung (2013), $\langle\delta_{r}u(\delta_{r}\theta)^{2}\rangle=-(4/3)\chi r$, holds in the scaling range. Here $\delta_{r}u$ is the increment of the velocity-like object. The factor $-4/3$ is for the three spatial dimensions. Testing against the Yaglom law is a reliable examination of the consistency, for it is an exact relationship for the passive scalar turbulence in the inertial-convective range. Unfortunately we do not have such a velocity-like object.
As shown in Fig. 8(b), the log of $|S_{3}(r)|$ suggests power-law behavior. Then we calculate the the third-order moments of the absolute value of the luminance, $\check{S}_{3}(r)=\langle|\delta_{r}\theta|^{3}\rangle$, which is shown in Fig. 8(c). Thanks to the absolute value, the curves become much less wiggly and the scaling exponents of $\check{S}_{3}(r)$ is close to the corresponding exponents, 0.873, of the passive scalar turbulence Watanabe and Gotoh (2004), adding another piece of consistency of the painting to the turbulence.
Refer to captionRefer to captionRefer to caption
Figure 8: Third-order structure function $S_{3}(r)$ in the linear-log coordinates (a) and its absolute value in the log-log coordinates (b). Third-order structure function of the absolute value of the luminance increment $\check{S}_{3}(r)=\langle|\delta_{r}\theta|^{3}\rangle$ is also shown in (c). The exponent $0.873$ in the panel (c) is taken from Ref.Watanabe and Gotoh, 2004 for the third-order moments of the absolute value of the passive scalar increment.
Note added in proof
A referee kindly pointed out the reference in vision science, D.J. Tolhurst, Y. Tadmor, and T. Chao, "Amplitude spectra of natural images", Ophthalmic and Physiological Optics 12, 229–232 (1992). In this reference, the luminance spectra of 135 photographs of diverse objects such as standing two persons, horses, natural scenery, a house and so forth, were analyzed. It is shown that their luminance spectra exhibit power-law scaling with the exponents ranging from $-2$ to $-0.6$, which include $-5/3$. This result demonstrates that agreement of the the power-law exponents of the spectra between a given image and the scalar-variance spectrum of turbulence does not necessarily imply a fundamental connection of the image to the passive scalar turbulence. In the present paper, we showed the agreement of not only the scaling of the luminance spectrum, but also the scaling of the structure function. Whether the structure functions of natural images exhibit a similar intermittent scaling law of the turbulent flow is an interesting subject to study.
[…]